# Learning incomplete factorization preconditioners for GMRES

### v1

@@ B.3 Baselines
The GMRES algorithm and the Arnoldi method, given in Algorithm 1 and Algorithm 2, respectively, are implemented in pytorch directly.
The Jacobi preconditioner can be implemented and applied very efficiently using pytorch sparse data structures for diagonal matrices. For the ILU(0) preconditioner, we use the ILU++ implementation which is a high-performance implementation of the algorithm in C++ (mayer2007ilupp; ilupp).
During inference, we run all operations – including the neural network-driven preconditioner – on the CPU in order to ensure a fair comparison between the different methods. For all methods, the time it takes to compute the preconditioner can be neglected compared to the time required for the GMRES algorithm.
@@ Appendix C Proofs
In the following section, we prove the lemmas from the main text. The proofs can easily be verified and rely on basic norm transformations and inequalities.
Proof of Lemma 1 In the first lemma, we are interested in obtaining an upper bound on the largest singular value of the preconditioned linear system. This allows our loss function to minimize an upper bound on the large end of the spectrum of the preconditioned system.
@@
$\leq\|{\bm{A}}-{\bm{P}}\|_{2}\;\|{\bm{P}}^{-1}\|_{2}+1$
Now, we exploit the structure of the learned preconditioner to obtain an upper bound for $\|{\bm{P}}^{-1}\|_{2}$. Specifically, we use two key observations. First, the matrix ${\bm{U}}$ has ones on the diagonal. Second, the absolute values of the diagonal elements on the matrix ${\bm{L}}$ are bounded from below by $\epsilon$, a property that arises from the specific choice of activation function in Equation (3).
@@
Proof of Lemma 2 In the second Lemma, we show how to obtain a lower bound on the smallest singular value of the preconditioned system. Our goal is to maximize this bound as it allows us to increase the small singular values of the preconditioned system.
@@
The reason for this is that optimizing first approximation directly leads to degenerate solutions as the minimum is attained when ${\bm{P}}\approx 0$ which would avoid having small singular values but the large singular value of the preconditioned system would be unbounded. Therefore, one can see the additional term involving ${\bm{I}}$ as a regularizer enforcing ${\bm{P}}$ to be non-degenerate.
Frobenius norm approximation  Equivalently to optimizing over the Frobenius norm, we can optimize the preconditioner over the squared Frobenius norm instead as we are only interested in the best parameters, not the best objective value. This has the additional advantage that the loss function is fully differentiable. Further, we can approximate the squared Frobenius norm using Hutchinson’s trace estimator which can be easily verified (hutchinson1989stochastic). For a random vector ${\bm{w}}$ that satisfies $\mathbb{E}\;[{\bm{w}}{\bm{w}}^{\mkern-1.5mu\mathsf{T}}]={\bm{I}}$ we can approximate the squared Froebnius norm using the following transformations:
@@
While it is possible to use multiple samples to obtain a more accurate approximation of the Frobenius norm, we follow previous approaches by estimating the norm using a single random vector to limit computational resources.
@@ Appendix D Additional results
Here, we provide additional results for the baseline and learned preconditioners. In Figure 5, we show the singular value distribution of the baseline preconditioners for the same problem as previously shown in Figure 3.
We can see that even though the ILU preconditioner has a few very large singular values, most values are clustered around $1$. This leads to a very fast convergence as observed in Table 1 even though the condition number $\kappa$ of the problem is very high.
Figure 5: Distribution of singular values of the preconditioned system for different choices of baseline preconditioner.

### v2

@@ B.3 Baselines
The GMRES algorithm and the Arnoldi method, given in Algorithm 1 and Algorithm 2, respectively, are implemented in pytorch directly. We utilize an iterative implementation of the Arnoldi method that only computes one orthogonal vector in each iteration for computational efficiency. As the considered problems here are of moderate size, we apply a direct QR factorization to the matrix ${\bm{H}}$ at each step, rather than incrementally updating the factorization through Givens rotations. Although Givens rotations are beneficial for larger or sparser systems due to their ability to maintain efficiency in incremental updates, a direct QR factorization is computationally feasible here and simplifies implementation golub2013matrix.
Classical Preconditioner  The Jacobi preconditioner can be implemented and applied very efficiently using pytorch sparse data structures for diagonal matrices. For the ILU(0) preconditioner, we use the ILU++ implementation which is a high-performance implementation of the algorithm in C++ (mayer2007ilupp; ilupp).
Data-driven preconditioner  The data-driven baseline uses an architecture similar to the one proposed by li2023learning. The architecture by hausner2023neural is not directly applicable since it only takes the lower triangular part of the matrix as an input which would remove many input features as the matrices in our dataset are non-symmetric. However, we use a $\exp$-activation function for the diagonal elements instead of choosing the original elements in ${\bm{A}}$ since the matrices in our dataset are not guaranteed to have non-zero diagonal elements. Further, we use the same node and edge features as our GNN described in the previous section.
Inference  During inference, we run all operations – including the neural network-driven preconditioner – on the CPU in order to ensure a fair comparison between the different methods. For all methods, the time it takes to compute the preconditioner can be neglected compared to the GMRES time.
@@ Appendix C Proofs
In this section, we prove the lemmas from the main text. The goal of Lemmas 1 and 2 is to optimize the system’s spectral properties by minimizing the upper and maximizing the lower spectral bounds respectively. This approach allows our loss function to increase the smaller singular values of the preconditioned system while decreasing the larger singular values, thereby improving the clustering of the singular values in the preconditioned matrix. The proofs can easily be verified and rely on basic norm transformations and norm inequalities.
Proof of Lemma 1 In the first lemma, we aim to upper bound the largest singular value of the preconditioned linear system. The bound is derived as follows:
@@
$\leq\|{\bm{A}}-{\bm{P}}\|_{2}\;\|{\bm{P}}^{-1}\|_{2}+1.$
Now, we exploit the structure of the learned preconditioner to obtain an upper bound for $\|{\bm{P}}^{-1}\|_{2}$. Specifically, we use two key observations. First, the matrix ${\bm{U}}$ has ones on the diagonal by construction. Second, the absolute values of the diagonal elements on the matrix ${\bm{L}}$ are bounded from below by $\epsilon$, a property that arises from the specific choice of activation function in Equation (3).
@@
Proof of Lemma 2 In the second lemma, we show how to obtain a lower bound on the smallest singular value of the preconditioned system and it is derived as follows:
@@
The reason for working with this second weaker approximation is that optimizing the first approximation directly leads to degenerate solutions as the minimum is attained when ${\bm{P}}\approx 0$ which would avoid having small singular values but the large singular value of the preconditioned system would be unbounded. Therefore, one can see the additional term involving ${\bm{I}}$ as a regularizer enforcing ${\bm{P}}$ to be non-degenerate.
Frobenius norm approximation  Equivalently to optimizing over the Frobenius norm, we can optimize the preconditioner over the squared Frobenius norm instead as we are only interested in the best parameters, not the best objective value. This has the additional advantage that the loss function is fully differentiable. Further, we can approximate the squared Frobenius norm using Hutchinson’s trace estimator which can be easily verified (hutchinson1989stochastic). For a random vector ${\bm{w}}$ that satisfies $\mathbb{E}\;[{\bm{w}}{\bm{w}}^{\mkern-1.5mu\mathsf{T}}]={\bm{I}}$ we can approximate the squared Frobenius norm using the following transformations:
@@
While it is possible to use multiple samples to obtain a more accurate approximation of the Frobenius norm, we follow previous approaches by estimating the norm using a single random vector to limit computational resources to create the dataset.
@@ Appendix D Additional results
Here, we present additional results for the singular values, choice of activation function, and robustness with respect to different hyper-parameters $\epsilon$.
D.1 Singular value distribution
Figure 5: Descending log-scale plot of ordered singular values of single problem instance for different baseline preconditioners.
We provide additional results for the baseline and learned preconditioners. In Figure 5 we present the descending log-scale plot of ordered singular values for the baseline preconditioners, illustrating the same problem previously depicted in Figure 3. We can see that even though the ILU preconditioner has a few very large singular values, most values are clustered around $1$ (Observe that the plot shows the logarithmic singular values thus the cluster appears at $0=\log 1$). This leads to a very fast convergence as observed in Table 1 even though the condition number $\kappa$ of the problem is very high.
This fact gets even clearer when taking a look at the distribution of singular values shown in Figure 6 for both the learned preconditioner and the baseline methods.
Figure 6: Density plot showing the distribution of singular values for a single problem instance across various preconditioners.
We can see that the Jacobi preconditioner does not change the distribution of singular values significantly but only shifts the spectrum towards $1$. Therefore, no improved convergence can be observed. The incomplete LU method leads to a tighter clustering of the singular values with a fast decay towards small singular values while some large singular values remain. For the learned methods, we can see that using the loss $\mathcal{L}_{\min}$ from Equation (6) leads to a distribution far away from zero but the singular values are not very clustered. The best clustering is achieved using the loss corresponding to large singular values $\mathcal{L}_{\max}$. However, the spectrum does not show a fast decay towards small singular values. In the spectrum of the combined loss, we can see the combined effect of the two approaches: the spectrum is clustered around a single value and the singular values are away from zero. These two properties of the spectrum make the preconditioned system easier to solve leading to better performance observed in the experiments presented in Table 1 in the main text.
D.2 Activation function
In order to evaluate the chosen activation function to ensure invertability, we compare the performance of the learned LU preconditioner with different values of the hyper-parameter $\epsilon$. Note that, the performance of the loss function and the chosen hyper-parameter also highly depends on the initial singular value distribution which is shifted to small singular values in our experiments as discussed previously.
The results of the learned preconditioner for $4$ different values of $\epsilon$ are shown in Table 2. We use the combined loss function $\mathcal{L}_{\text{com}}$ for all experiments. We can see that with a larger $\epsilon$ parameter, both the small singular values of the preconditioned system as well as the large singular values decrease. The best performance is achieved when balancing the two different ends of the singular value spectrum.
Finally, we can see that all models trained with the combined loss outperform the previous models trained on only one loss function showing the robustness of our proposed method.
Table 2: Performance of our method for different values of $\epsilon$ controlling the activation function. The columns presented here are a subset of the columns shown in Table 1.
$\epsilon$ $\sigma_{\min}$uparrow $\sigma_{\max}$downarrow $\kappa$downarrow Its. $\downarrow$
0.0001 0.0022 39.12 37 513 420
0.001 0.0015 35.17 49 723 418
0.01 0.0012 17.26 30 349 428
0.1 0.0010 12.39 24 702 424
Connection to other data-driven techniques  Noteworthy, the method proposed by li2023learning employs the loss function $\mathcal{L}_{\min}$, which leads to better results compared to those observed in our experiments. However, a crucial difference between this method and our approach is the fact that the diagonal elements are not learned in the NeuralPCG preconditioner. Instead, the square root of the original elements of the matrix ${\bm{A}}$ is used. Since the matrices considered are positive definite, the diagonal elements are non-zero, and therefore, the incomplete factorization is guaranteed to be invertible. For general invertible matrices, that we explore in this paper, this is, however, not the case. The theoretical results presented here explain some of the discrepancies between our results and the previously obtained improvements. Implicitly, li2023learning choose a large $\epsilon$ for these matrices by construction of the diagonal elements. This allows them to bound the largest singular value under the assumption that the difference between the original matrix and the learned factorization is not too large by applying Lemma 1. Further, it is easy to integrate the method developed here with data-driven methods for incomplete Cholesky methods (hausner2023neural). This can be achieved by adding the $\epsilon$ parameter directly to the diagonal elements in order to ensure a lower bound on the values.